# Disentangled Representation Learning through Geometry Preservation with the Gromov-Monge Gap

### v1

@@ 3 Disentanglement with the Gromov-Monge gap
This section details our novel optimal transport perspective to achieve disentanglement from geometric considerations (see § 2.1), using the VAE framework. To achieve this, we first investigate how one can promote an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$ between two domains $\mathcal{X}$ and $\mathcal{Y}$ to preserve predefined geometric features. In a VAE, $T$ can represent either the encoder $e_{\phi}$, which produces latent codes from the data, or the decoder $d_{\theta}$, which reconstructs the data from the latent codes. As a result, in the former, the source domain $\mathcal{X}$ is the data, and the target domain $\mathcal{Y}$ is the latent space, with roles swapped in the latter. If we assume that $d_{\theta}$ perfectly reconstructs the data from the latents produced by $e_{\phi}$, it is equivalent whether $e_{\phi}$ preserves the geometric features from data to latents or $d_{\theta}$ preserves them from latents to data. Consequently, in what follows, $T$ can refer to either the encoder or the decoder without distinction.
@@
Leveraging this perspective, this section begins by defining cost functions to encode geometric features and the notion of distortions in §3.1. We leverage this concept in §3.2 to introduce the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves points while preserving geometric features as much as possible, i.e., minimizing distortion. §3.3 then shows how the GMG can be estimated and computed from samples to be practically applicable in the VAE framework, which transitions into §3.4 studying convexity properties of the GMG. Put together, §3.13.4 define the practical GMG which allows us to learn a latent space that matches, as much as possible, geometrical constraints in the data space. Finally in §3.5, we use the GMG with different choices of costs to propose effective disentangled representation learning objectives.
@@ 3.1 From the distortion…
We encode the geometric features of interest through two cost functions defined on each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. We then want $T$ to preserve these costs, i.e., $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}$. In practice, two types of cost functions are particularly meaningful.
@@
    (Scaled) squared Euclidean distance: $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}$ and $c_{\mathcal{Y}}=\alpha^{2}\|\mathbf{y}-\mathbf{y}\|_{2}^{2}$, with $\alpha>0$. A map $T$ preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ preserves the scaled distances between the points, i.e. it is a scaled isometry. When $\alpha=1$, we recover the standard definition of an isometry.
@@
In the following, we say that $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] if they belong to these families of costs. Introducing a reference distribution $r\in\mathcal{P}(\mathcal{X})$, weighting the areas of $\mathcal{X}$ where we penalize deviations of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ from $c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$, we can quantify this property using the following criterion.
@@
The distortion (DST) of a map $T$ is defined as:
$\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T):=\int_{\mathcal{X}\times\mathcal{X}}\tfrac{1}{2}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})$ (DST)
$\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ quantifies how much $T$ distorts the geometric features induced by $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ on the support of $r$, i.e., when $\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=0$, one has $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathrm{Spt}(r)$. In disentangled representation learning, it can be desirable to regularize the decoder to be isometric [29, 51, 31]. However, a fully geometry-preserving mapping might not necessarily exist between the latent space and the data distribution. This means there usually exists an inherent trade-off between the accurate reconstruction of the data distribution and this reconstruction being, e.g., a "fully isometric" map. If such a map does not exist, the reconstruction loss and the distortion term cannot be 00 simultaneously. In practice, this means that the distortion loss will move away from accurately reconstructing the data, which negatively impacts the quality of the learned latent representations. This naturally raises the question of how to formulate a geometric regularization that takes the most geometry-preserving mapping into account.
@@
3.2 …to The Gromov-Monge Gap
Recently, Uscidda and Cuturi [68] introduced the Monge gap, a regularizer that measures whether a map $T$ transports a reference distribution at the minimal displacement cost. In practice, this regularizer is combined with fitting losses to compute Monge maps, which are defined by two main features: (i) they fit a marginal constraint with (ii) minimal displacement cost. Building on this concept, we replace "displacement" with "distortion" to introduce the Gromov-Monge gap, a regularizer that assesses whether a map $T$ transports a reference distribution at the minimal distortion cost. In § 3.5, we use it, alongside fitting losses, to compute Gromov-Monge maps, as defined in Eq. (GMP), which are similarly defined by (i) fitting a marginal constraint with (ii) minimal distortion cost.
Definition 3.2 (Gromov-Monge gap).
The Gromov-Monge gap (GMG) of a map $T$ is defined as:
$\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\vcentcolon=\mathcal{D}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)$ (GMG)
From Eq. (GWP), we recall that $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)$ is the minimal distortion achievable when transporting $r$ to $T\sharp r$. Thus, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ quantifies the difference between the distortion incurred when transporting $r$ to $T\sharp r$ via $T$, to this minimal distortion. More formally, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ is the optimality gap of $T$ in the Gromov-Monge Prob. (GMP) between $r$ and $T\sharp r$, which is always feasible, even when $r$ is discrete, as $T$ belongs to the constraint set. In light of this, it is a well-defined and
  • The GMG measures how close $T$ is to be a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. Indeed, $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\geq 0$ with equality i.f.f. $T$ is a Gromov-Monge map solution of Prob. (GMP) between $r$ and $T\sharp r$, i.e., $T$ moves $r$ with minimal (but eventually non zero) distortion.
  • When transport without distortion is possible, the GMG coincides with the distortion. When there exists another map $U:\mathcal{X}\to\mathcal{Y}$ transporting $r$ to $T\sharp r$ with zero distortion, i.e., $U\sharp r=T\sharp r$ and $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(U)=0$, then $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)$. Indeed, $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r,T\sharp r)=0$ in that case, as the coupling $\pi=(\mathrm{Id},U)\sharp r$ sets the GW objective to zero, thereby minimizing it.
The last point (ii) is fundamental and illustrates how the GMG functions as a debiased distortion. Indeed, it compares the distortion induced by $T$ to a baseline distortion, defined as the minimal achievable distortion when transforming the reference distribution into its image under $T$. Thus, when transformation without any distortion is achievable, the reference distortion becomes zero, and the GMG precisely aligns with the distortion itself, i.e., $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)$. In this context, the GMG offers the optimal compromise: it avoids the over-penalization induced by the distortion when fully preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ is not feasible, yet it coincides with it when such preservation is feasible.
The Influence of the Reference Distribution.
A crucial property of $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}$ is that if $T$ transforms $r$ without distortion, it will also apply distortion-free to any distribution $s$ whose support is contained within that of $r$. Formally, if $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)=0$ and $s\in\mathcal{P}(\mathcal{X})$ with $\operatorname{supp}(s)\subseteq\operatorname{supp}(r)$, then $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{s}(T)=0$. This raises a question for the GMG: If $T$ maps $r$ with minimal distortion, does it similarly map $s$ with minimal distortion? We answer this question positively with Prop. (3.3) when the costs are the (scaled) Euclidean distances or the cosine similarity. Intuitively, this means that if $T$ moves $r$ while preserving (scaled) distances or angles as much as possible, it will also preserve these properties as much as possible when moving any "smaller" distribution within $r$.
Proposition 3.3.
When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] (see § 3.1), if $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}(T)=0$, then for any $s\in\mathcal{P}(\mathcal{X})$ s.t. $\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, one has $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{s}(T)=0$.
3.3 Estimation and Computation from Samples
Plug-In Estimation.
In practice, we estimate Eq. (DST) and Eq. (GMG) using i.i.d. samples $\mathbf{x}_{1},...,\mathbf{x}_{n}$ from the reference distribution $r$. We then consider the empirical version ${r}_{n}\vcentcolon=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ of $r$ and use a plug-in estimator for both cases, i.e., we estimate the distortion via
$\begin{split}\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{i}),T(\mathbf{x}_{j})))^{2}\,,\end{split}$ (2)
and the GMG via $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n}}(T)=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$, where $T\sharp{r}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$. To better understand what the discrete GMG quantifies, we can reformulate it using the minimal distortion achieved by a permutation $\sigma\in\mathcal{S}_{n}$ between the $\mathbf{x}_{i}$ and the $T(\mathbf{x}_{i})$.
Proposition 3.4.
When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii] , the empirical GMG reads:
$\begin{split}\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\min_{\sigma\in\mathcal{S}_{n}}\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\left(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{\sigma(i)}),T(\mathbf{x}_{\sigma(j)}))\right)^{2}\end{split}$ (3)
As a Monte Carlo estimator, $\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ is naturally consistent. We can ask the same question for $\mathcal{GM}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n}}(T)$, which requires studying the convergence of the empirical GW distance $\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$. For the costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ of interest, we show that consistency holds.
Proposition 3.5.
When $c_{\mathcal{X}},c_{\mathcal{Y}}$ are [i] or [ii], $\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\rightarrow\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ almost surely.
Efficient Computation.

### v2

@@ 3 Disentanglement with the Gromov-Monge gap
This section outlines our novel OT-based approach to disentanglement through geometric considerations within the VAE framework. First, we explore how an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$ between two domains can be encouraged to preserve predefined geometric features. In a VAE, $T$ may represent either the encoder $e_{\phi}$, which generates latent codes from the data, or the decoder $d_{\theta}$, which reconstructs the data from these codes. In the case of the encoder, the source domain $\mathcal{X}$ corresponds to the data, and the target domain $\mathcal{Y}$ is the latent space, while for the decoder, these roles are swapped. Assuming that $d_{\theta}$ perfectly reconstructs the data from the latents produced by $e_{\phi}$, i.e., $e_{\phi}\circ d_{\theta}=\mathrm{I}d$, the preservation of geometric features by either $e_{\phi}$ or $d_{\theta}$ becomes equivalent. Therefore, in the following sections, we refer to $T$ as either the encoder or the decoder without loss of generality.
@@
Building on this perspective, this section begins by defining cost functions to encode geometric features and the notion of distortions in §3.1. We leverage this concept in §3.2 to introduce the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves distributions while preserving geometric features as much as possible, i.e., minimizing distortion. §3.3 then shows how the GMG can be efficiently computed from samples to be practically applicable in the VAE framework. This transitions into §3.4 studying (weak) convexity properties of the GMG, as an operator. Finally, in §3.5, we describe how to integrate the GMG with disentangled representation learning objectives, effectively combining prior matching with geometric constraints.
@@ 3.1 From the distortion…
We encode the geometric features of interest using two cost functions, one for each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. We aim for $T$ to preserve the geometric feature, which means that $T$ preserves these costs, that is, $c_{\mathcal{X}}(\mathbf{x},\mathbf{x^{\prime}})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x^{\prime}}))$ for $\mathbf{x},\mathbf{x^{\prime}}\in\mathcal{X}$. In practice, two types of cost function from this family are particularly relevant:
@@
    (Scaled) squared L2 distance: $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})=\alpha^{2}\|\mathbf{y}-\mathbf{y}\|_{2}^{2}$, with $\alpha>0$. A map $T$ preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ preserves the scaled distances between the points, i.e. it is a scaled isometry. When $\alpha=1$, we recover the standard definition of an isometry.
@@
We refer to these costs via $\mathbf{L2^{2}}$ for [i] with $\alpha=1$, $\mathbf{ScL2^{2}}$ for [i] with $\alpha\neq 1$ and $\mathbf{Cos}$ for [ii]. Introducing a reference distribution $r\in\mathcal{P}(\mathcal{X})$, weighting the areas of $\mathcal{X}$ where we penalize deviations of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ from $c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$, we can quantify this property using the following criterion:
@@
The distortion (DST) of a map $T$, for cost functions $c_{\mathcal{X}},c_{\mathcal{Y}}$ and reference distribution $r$, is defined as:
$\mathrm{DST}_{r}(T):=\int_{\mathcal{X}\times\mathcal{X}}\tfrac{1}{2}(c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime})))^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})\mathop{}\!\mathrm{d}r(\mathbf{x}^{\prime})\,.$ (DST)
$\mathrm{DST}_{r}(T)$ quantifies how much $T$ distorts geometric features encoded by $c_{\mathcal{X}},c_{\mathcal{Y}}$ on the support of $r$, that is, when $\mathrm{DST}_{r}(T)=0$, one has $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathrm{Spt}(r)$. Nakagawa et al. (2023) suggest promoting geometry preservation by regularizing the encoder $d_{\theta}$ using the DST, with $\mathbf{ScL2^{2}}$ as costs, and the latent representation as reference distribution, namely $r=e_{\phi}\sharp p_{\textrm{data}}$. While they use it within a WAE, translated in our VAE setting, this corresponds to minimizing
$\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}]+\underbrace{\lambda\,\mathrm{DST}_{r}(d_{\theta})}_{\textrm{(iii) geom. preservation}},$ (2)
where $\lambda>0$. Given the choice of the costs, $d_{\theta}$ is distortion-free (i.e., $\mathrm{DST}_{r}(d_{\theta})=0$), if it is a scaled isometry. Since a scaled isometry that maps the prior onto the data distribution may not exist, there is an inherent trade-off between minimizing terms (ii) and (iii), which are responsible for achieving practical disentanglement. As these terms cannot be simultaneously minimized to 00, the DST loss will move away from accurately matching the prior, which will negatively impact the quality of the learned latent representations. This naturally raises the question of how to avoid this over-penalization. Instead of seeking a distortion-free decoder, we should seek a decoder that transports the prior to the data distribution with minimal distortion of the costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. In other words, the decoder should be a Gromov-Monge map between the prior and the data distributions for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ (see 2.2). Conversely, if we choose to regularize the encoder, the same reasoning applies by swapping the roles of the prior and the data distribution. In the next section, we introduce a regularizer to fit Gromov-Monge maps, which we will use as a replacement to the DST.
3.2 …to The Gromov-Monge Gap
Figure 1: Sketch of $\mathrm{GMG}_{\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}}(\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}})$ for two different maps $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$. We use a discrete reference distribution $\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}$ on 3 points, and $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}$, hence measuring if $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$ minimally distorts the distances. On the left, $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$ is the optimal map, $\boldsymbol{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T^{\star}}}$, and maps the three points with minimal (yet non-zero) distortion, which is measured as the sum of the squared lengths of the orange segments. This results in $\mathrm{GMG}_{\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}}(\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}})=0$. On the right, $\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}}$ swaps two points compared to $\boldsymbol{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T^{\star}}}$—specifically, $\boldsymbol{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}_{2}}}$ and $\boldsymbol{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}_{3}}}$—causing a higher distortion than the minimal one, and measured as the sum of the squared lengths of the red segments. This results in $\mathrm{GMG}_{\boldsymbol{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}r_{n}}}}(\boldsymbol{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}})>0$, equal to the gap between these distortions.
Recently, Uscidda and Cuturi (2023) introduced the Monge gap, a regularizer that measures whether a map $T$ transports a reference distribution at the minimal displacement cost. Building on this concept, we replace "displacement" with "distortion" to introduce the Gromov-Monge gap, a regularizer that assesses whether a map $T$ transports a reference distribution at the minimal distortion cost.
Definition 3.2 (Gromov-Monge gap).
The Gromov-Monge gap (GMG) of a map $T$, for cost functions $c_{\mathcal{X}},c_{\mathcal{Y}}$ and reference distribution $r$, is defined as:
$\mathrm{GMG}_{r}(T)\vcentcolon=\mathrm{DST}_{r}(T)-\mathrm{GW}(r,T\sharp r)$ (GMG)
We recall from Eq. (GWP) that $\mathrm{GW}(r,T\sharp r)$ represents the minimal distortion of $c_{\mathcal{X}},c_{\mathcal{Y}}$ achievable when transporting $r$ to $T\sharp r$. Thus, the GMG quantifies the difference between the distortion incurred when transporting $r$ to $T\sharp r$ via $T$, and this minimal distortion. Formally, when Prob. (GMP) and Prob. (GWP) between $r$ and $T\sharp r$ are equivalent, the the GMG is the suboptimality gap of $T$ in Prob. (GMP). This is the case, for example, when $r$ is a density and $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ (Dumont et al., 2022). Otherwise, the GMG is the suboptimality gap of $\pi=(\mathrm{I}_{d},T)\sharp r$ in Prob. (GWP) between $r$ and $T\sharp r$. In light of this, it is a well-defined quantity and:
  • The GMG measures how close $T$ is to be a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. Indeed, $\mathrm{GMG}_{r}(T)\geq 0$ with equality if $T$ is a Gromov-Monge map solution of Prob. (GMP) between $r$ and $T\sharp r$, i.e., $T$ moves $r$ with minimal (but eventually non zero) distortion; see App. B.1.
  • When transport without distortion is possible, the GMG coincides with the distortion. When there exists another map $U:\mathcal{X}\to\mathcal{Y}$ transporting $r$ to $T\sharp r$ with zero distortion, i.e., $U\sharp r=T\sharp r$ and $\mathrm{DST}_{r}(U)=0$, then $\mathrm{GMG}_{r}(T)=\mathrm{DST}_{r}(T)$. Indeed, $\mathrm{GW}(r,T\sharp r)=0$ in that case, as the coupling $\pi=(\mathrm{Id},U)\sharp r$ sets the GW objective to zero, thereby minimizing it.
The last point illustrates how the GMG functions as a debiased distortion. It compares the distortion induced by $T$ to a baseline distortion, defined as the minimal achievable distortion when transforming $r$ into $T\sharp r$. Thus, when transformation without distortion is achievable, this baseline becomes zero, and the GMG equals the distortion. Consequently, the GMG offers the optimal compromise: it avoids the over-penalization induced by the distortion when fully preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ is not feasible, yet it coincides with it when such full preservation is feasible. See Fig. 1 for a simple illustration.
The Influence of the Reference Distribution.
A crucial property of $\mathrm{DST}_{r}$ is that if $T$ transforms $r$ without distortion, it will also apply distortion-free to any distribution $s$ whose support is contained within that of $r$. Formally, if $\mathrm{DST}_{r}(T)=0$ and $s\in\mathcal{P}(\mathcal{X})$ with $\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, then $\mathrm{DST}_{s}(T)=0$. This raises a question for the GMG: If $T$ maps $r$ with minimal distortion, does it similarly map $s$ with minimal distortion? We answer this question positively with Prop. (3.3).
Proposition 3.3.
If $\mathrm{GMG}_{r}(T)=0$, $\forall s\in\mathcal{P}(\mathcal{X})$ s.t. $\mathrm{Spt}(s)\subseteq\mathrm{Spt}(r)$, one has $\mathrm{GMG}_{s}(T)=0$.
3.3 Estimation and Computation from Samples
Plug-In Estimation. In practice, we estimate Eq. (DST) and Eq. (GMG) from samples $\mathbf{x}_{1},...,\mathbf{x}_{n}\sim r$. We consider the empirical version ${r}_{n}\vcentcolon=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ of $r$ and use plug-in estimators, i.e.
$\begin{split}\mathrm{DST}_{{r}_{n}}(T)=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(T(\mathbf{x}_{i}),T(\mathbf{x}_{j})))^{2}\,,\end{split}$ (3)
and $\mathrm{GMG}_{{r}_{n}}(T)=\mathrm{DST}_{{r}_{n}}(T)-\mathrm{GW}({r}_{n},T\sharp{r}_{n})$, where $T\sharp{r}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$.
Algorithm 1 GMG$(\mathbf{x}_{1},\dots\mathbf{x}_{n},T,\varepsilon)$.
@@ 3.3 Estimation and Computation from Samples
2:$\mathbf{t}_{1},\dots\mathbf{t}_{n}\leftarrow T(\mathbf{x}_{1}),\dots,T(\mathbf{x}_{n})$.
3:$\mathbf{C}_{\mathcal{X}}\leftarrow[c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})]_{1\leq i,i^{\prime}\leq n}$    $\triangleright$ usually $\mathcal{O}(n^{2}d_{\mathcal{X}})$
4:$\mathbf{C}_{\mathcal{Y}}\leftarrow[c_{\mathcal{Y}}(\mathbf{t}_{j},\mathbf{t}_{j^{\prime}})]_{1\leq j,j^{\prime}\leq n}$    $\triangleright$ usually $\mathcal{O}(n^{2}d_{\mathcal{Y}}$)
5:$\textsc{DST}\leftarrow\texttt{mean}((\mathbf{C}_{\mathcal{X}}-\mathbf{C}_{\mathcal{Y}})^{2})$triangleright $n^{2}$
6:$\varepsilon\leftarrow\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$triangleright usually $\mathcal{O}(n^{2})$
7:$\textsc{GW}\leftarrow\texttt{GW}(\mathbf{C}_{\mathcal{X}},\mathbf{C}_{\mathcal{Y}},\varepsilon)$triangleright $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$
8:return DST - GW
Efficient and Stable Computation.